Which of the above equilibria is forced to shift furthest to the left? Explain.

- Lead nitrate and Sodium chloride
- Lead chloride and Sodium chromate
-Lead chromate and Sodium sulphide
1
Expert's answer
2020-03-04T07:59:54-0500

Answer: Combination which makes less soluble product shift towards left side of the equation.

Pb(NO3)2 + 2NaCl = PbCl2 + 2NaNO3

PbCl2 + Na2CrO4 = PbCrO4 + 2NaCl

PbCrO4 + Na2S = PbS + 2NaCrO4

PbS and PbCrO4 is less soluble in water, so equilibrium will shift towards left side in this two cases.